OpenStudy (alesia/bailey):

Please help, I need someone to recheck this. 15x+24y/-3

OpenStudy (alesia/bailey):

I ended up getting 15x+24y

jimthompson5910 (jim_thompson5910):

The original expression is \[\Large \frac{15x+24y}{-3}\] right?

OpenStudy (alesia/bailey):

yes

jimthompson5910 (jim_thompson5910):

So what you do is break up the fraction to get \[\Large \frac{15x+24y}{-3}\] \[\Large \frac{15x}{-3}+\frac{24y}{-3}\] then you divide 15 by -3 to get -5. Also you divide 24 by -3 to get -8 so \[\Large \frac{15x}{-3}+\frac{24y}{-3}\] turns into \[\Large -5x - 8y\]

jimthompson5910 (jim_thompson5910):

Hopefully that makes sense?

jimthompson5910 (jim_thompson5910):

or are you stuck somewhere?